# Subsampled Ensemble Can Improve Generalization Tail Exponentially

### v3

@@ 4 Related work
This work is closely connected to various topics in optimization and machine learning, and we only review the most relevant ones. See Appendix A for additional literature review.
Ensemble learning.
@@
Ensemble learning (Dietterich, 2000; Zhou, 2012; Sagi & Rokach, 2018) has been widely studied for improving model performance by combining multiple weak learners into strong ones. Popular ensemble methods include bagging (Breiman, 1996), boosting (Freund et al., 1996) and stacking (Wolpert, 1992; Džeroski & Ženko, 2004). Bagging enhances model stability by training models on different bootstrap samples and combining their predictions through majority voting or averaging, effectively reducing variance, especially for unstable learners like decision trees that underpin random forests (Breiman, 2001). Subagging (Bühlmann & Yu, 2002) is a variant of bagging that constructs the ensemble from subsamples in place of bootstrap samples. Boosting is a sequential process where each subsequent model corrects its predecessors’ errors, reducing both bias and variance (Ibragimov & Gusev, 2019; Ghosal & Hooker, 2020). Prominent boosting methods include AdaBoost (Freund et al., 2003), Stochastic Gradient Boosting (SGB) (Friedman, 2001; 2002), and Extreme Gradient Boosting (XGB) (Friedman et al., 2000) which differ in their approaches to weighting training data and hypotheses. Boosting is commonly used with decision trees as Gradient Boosted Decision Trees (GBDT), including XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Hancock & Khoshgoftaar, 2020). Instead of using simple aggregation like weighted averaging or majority voting, stacking trains a model to combine base predictions in a more sophisticated way, further improving performance. A key procedural difference of our approach from these ensemble methods is that we perform majority voting at the model level, rather than at the prediction level, to select a single best model from the ensemble. As a result, our method consistently outputs models within the same space as the base learner, making it applicable to general stochastic optimization problems. In contrast, most existing ensemble methods yield aggregated models outside the base space. Additionally, compared to the bias/variance reduction of typical ensembles, our approach guarantees exponentially decaying excess risk tails and hence is particularly effective in settings with heavy-tailed noise.
Optimization and learning with heavy tails.
@@
Optimization with heavy-tailed noises has garnered significant attention due to its relevance in traditional fields such as portfolio management (Mainik et al., 2015) and scheduling (Im et al., 2015), as well as emerging domains like large language models (Brown et al., 2020; Achiam et al., 2023). Tail bounds of most existing algorithms are guaranteed to decay exponentially under sub-Gaussian or uniformly bounded costs but deteriorate to a slow polynomial decay under heavy-tailedness (Kaňková & Houda, 2015; Jiang et al., 2020; Jiang & Li, 2021; Oliveira & Thompson, 2023). For SAA or ERM, faster rates are possible under the small-ball (Mendelson, 2018; 2015; Roy et al., 2021) or Bernstein’s condition (Dinh et al., 2016) on the function class, while our approach is free from such conditions. Considerable effort has been made to mitigate the adverse effects of heavy-tailedness with robust procedures among which the geometric median (Minsker, 2015), or more generally, median-of-means (MOM) (Lugosi & Mendelson, 2019a; c) approach is most similar to ours. The basic idea there is to estimate a true mean by dividing the data into disjoint subsamples, computing an estimate on each, and then taking the median. Lecué & Lerasle (2019); Lugosi & Mendelson (2019b); Lecué & Lerasle (2020) use MOM in estimating the expected cost and establish exponential tail bounds for the mean squared loss and convex function classes. Hsu & Sabato (2016; 2014) apply MOM directly on the solution level for continuous problems and require strong convexity from the cost to establish generalization bounds. Besides MOM, another approach estimates the expected cost via truncation (Catoni, 2012) and allows heavy tails for linear regression (Audibert & Catoni, 2011; Zhang & Zhou, 2018) or problems with uniformly bounded function classes (Brownlees et al., 2015), but is computationally intractable due to the truncation and thus more of theoretical interest. In contrast, our ensemble approach is a meta algorithm that acts on any learning algorithm to provide exponential tail bounds regardless of the underlying problem characteristics. Relatedly, various techniques such as gradient clipping (Cutkosky & Mehta, 2021; Gorbunov et al., 2020) and MOM (Puchkin et al., 2024) have been adopted in stochastic gradient descent (SGD) algorithms for handling heavy-tailed gradient noises, but their focus is the faster convergence of SGD rather than generalization.
Machine learning for optimization.
Learning to optimize (L2O) studies the use of machine learning in accelerating existing or discovering novel optimization algorithms. Much effort has been in training models via supervised or reinforcement learning to make critical algorithmic decisions such as cut selection (e.g., Deza & Khalil (2023); Tang et al. (2020)), search strategies (e.g., Khalil et al. (2016); He et al. (2014); Scavuzzo et al. (2022)), scaling (Berthold & Hendel, 2021), and primal heuristics (Shen et al., 2021) in mixed-integer optimization, or even directly generate high-quality solutions (e.g., neural combinatorial optimization pioneered by Bello et al. (2016)). See Chen et al. (2022; 2024); Bengio et al. (2021); Zhang et al. (2023) for comprehensive surveys on L2O. This line of research is orthogonal to our goal, and L2O techniques can work as part of or directly serve as the base learning algorithm within our framework.
5 Conclusion and limitation
@@
This paper introduces a novel ensemble technique that significantly improves generalization by aggregating base learners via majority voting. In particular, our approach converts polynomially decaying generalization tails into exponential decay, thus providing order-of-magnitude improvements as opposed to constant factor improvements exhibited by variance reduction. Extensive numerical experiments in both machine learning and stochastic programming validate its effectiveness, especially for scenarios with heavy-tailed data and slow convergence rates. This work underscores the powerful potential of our new ensemble approach across a broad range of machine learning applications.
While our method accelerates tail convergence, it may increase model bias, similar to other subsampling-based techniques like subagging (Bühlmann & Yu, 2002). This makes it best suited for applications with relatively low bias, e.g., when the model is sufficiently expressive.
@@
The appendices are organized as follows. In Appendix A, we review additional related work. Appendix B presents additional technical discussion for Theorem 1. Next, in Appendix C, we document the proofs of theoretical results in our paper. Specifically, we introduce some preliminary definitions and lemmas in Appendix C.1. Then, the proof of Theorem 1 can be found in Appendix C.2, and the proof of Theorem 2 can be found in Appendix C.3. Finally, we provide additional numerical experiments in Appendix D.
@@ Appendix A Additional related work
Bagging has been adopted in stochastic optimization for various purposes. The most relevant line of works (Biggs et al., 2023; Perakis & Thayaparan, 2021; Wang et al., 2021; Biggs & Perakis, 2023) study mixed integer reformulations for stochastic optimization with bagging approximated objectives such as random forests and ensembles of neural networks with the ReLU activation. These works focus on computational tractability instead of generalization performance. Anderson & Nguyen (2020) empirically evaluates several statistical techniques including bagging against the plain SAA and finds bagging advantageous for portfolio optimization problems. Birge (2023) investigates a batch mean approach for continuous optimization that creates subsamples by dividing the data set into non-overlapping batches instead of resampling and aggregates SAA solutions on the subsamples via averaging, which is empirically demonstrated to reduce solution errors for constrained and high-dimensional problems. Another related batch of works (Lam & Qian, 2018a; b; Chen & Woodruff, 2024; 2023; Eichhorn & Römisch, 2007) concern the use of bagging for constructing confidence bounds for generalization errors of data-driven solutions, but they do not attempt to improve generalization. Related to bagging, bootstrap has been utilized to quantify algorithmic uncertainties for randomized algorithms such as randomized least-squares algorithms (Lopes et al., 2018), randomized Newton methods (Chen & Lopes, 2020), and stochastic gradient descent (Fang et al., 2018; Zhong et al., 2023), which is orthogonal to our focus on generalization performance.
Appendix B Implications of Theorem 1 for strong base learners
@@
We provide a brief discussion of Theorem 1 applied to fast convergent base learners. Based on Theorem 1, the way $p_{k}^{\max}$ and $\mathcal{E}_{k,\delta}$ enter into (9) reflects how the generalization performance of the base learning algorithm is inherited by our framework. To explain, large $p_{k}^{\max}$ and small $\mathcal{E}_{k,\delta}$ correspond to better generalization of the base learning algorithm. This can be exploited by the bound (9) with the presence of $\max(1-p_{k}^{\max},\;\mathcal{E}_{k,\delta})$, which is captured with our sharper concentration of U-statistics with binary kernels. In particular, for base learning algorithms with fast generalization convergence, say $1-p_{k}^{\max}=\mathcal{O}(e^{-k})$ and $\mathcal{E}_{k,\delta}=\mathcal{O}(e^{-k})$ for simplicity, we have $C_{1}\max(1-p_{k}^{\max},\;\mathcal{E}_{k,\delta})=\mathcal{O}(e^{-k})$ and hence the first term in (9) becomes $\mathcal{O}(e^{-n})$ which matches the error of the base learning algorithm applied directly to the full data set.
Appendix C Technical proofs
@@ C.1 Preliminaries
An important tool in the development of our theories is the U-statistic that naturally arises in subsampling without replacement. We first present the definition of U-statistic and its concentration properties.
Definition 1
@@
Lemma 1 (MGF dominance of U-statistics from Hoeffding (1963))
For any integer $0<k\leq n$ and any kernel $\kappa(z_{1},\ldots,z_{k})$, let $U(z_{1},\ldots,z_{n})$ be the corresponding U-statistic defined in Definition 1, and

### v4

@@ 4 Related Work
This work is closely connected to various topics in optimization and machine learning, and we only review the most relevant ones. See Appendix A for additional reviews.
@@
Ensemble Learning.    Ensemble learning (Dietterich, 2000; Zhou, 2012; Sagi & Rokach, 2018) improves model performance by combining multiple weak learners into strong ones. Popular ensemble methods include bagging (Breiman, 1996), boosting (Freund et al., 1996) and stacking (Wolpert, 1992; Džeroski & Ženko, 2004). Bagging enhances stability by training models on different bootstrap samples and combining their predictions through majority voting or averaging, effectively reducing variance, especially for unstable learners like decision trees that underpin random forests (Breiman, 2001). Subagging (Bühlmann & Yu, 2002) is a variant of bagging that constructs the ensemble from subsamples in place of bootstrap samples. Boosting is a sequential process where each subsequent model corrects its predecessors’ errors, reducing both bias and variance (Ibragimov & Gusev, 2019; Ghosal & Hooker, 2020). Prominent boosting methods include AdaBoost (Freund et al., 2003), Stochastic Gradient Boosting (SGB) (Friedman, 2001, 2002), and Extreme Gradient Boosting (XGB) (Friedman et al., 2000) which differ in their approaches to weighting training data and hypotheses. Instead of using simple aggregation like weighted averaging or majority voting, stacking trains a model to combine base predictions to further improve performance. A key procedural difference of our approach from these methods is that we perform majority voting at the model rather than prediction level to select a single best model from the ensemble. That is, our approach outputs models in the same space as the base learner, whereas existing ensemble methods yield aggregated models outside the base space. This also means a constant inference cost for our output model with respect to the ensemble size, as opposed to linearly growing costs seen in existing ensemble methods. Methodologically, our approach operates by accelerating excess risk tail convergence in lieu of bias/variance reduction, and hence is particularly effective in settings with heavy-tailed noise.
@@
Optimization and Learning With Heavy Tails.    Optimization with heavy-tailed noises has garnered significant attention due to its relevance in traditional fields such as portfolio management (Mainik et al., 2015) and scheduling (Im et al., 2015), as well as emerging domains like large language models (Brown et al., 2020; Achiam et al., 2023). Tail bounds of most existing algorithms are guaranteed to decay exponentially under sub-Gaussian or uniformly bounded costs but deteriorate to a slow polynomial decay under heavy-tailedness (Kaňková & Houda, 2015; Jiang et al., 2020; Jiang & Li, 2021; Oliveira & Thompson, 2023). For SAA or ERM, faster rates are possible under the small-ball (Mendelson, 2018, 2015; Roy et al., 2021) or Bernstein’s condition (Dinh et al., 2016) on the function class, while our approach is free from such conditions. Considerable effort has been made to mitigate the adverse effects of heavy-tailedness with robust procedures among which the geometric median (Minsker, 2015), or more generally, median-of-means (MOM) (Lugosi & Mendelson, 2019a, c) approach is most similar to ours. The basic idea there is to estimate a true mean by dividing the data into disjoint subsamples, computing an estimate on each, and then taking the median. (Lecué & Lerasle, 2019; Lugosi & Mendelson, 2019b; Lecué & Lerasle, 2020) use MOM in estimating the expected cost and establish exponential tail bounds for the mean squared loss and convex function classes. (Hsu & Sabato, 2016, 2014) apply MOM directly on the solution level for continuous problems and require strong convexity from the cost to establish generalization bounds. Besides MOM, another approach estimates the expected cost via truncation (Catoni, 2012) and allows heavy tails for linear regression (Audibert & Catoni, 2011; Zhang & Zhou, 2018) or problems with uniformly bounded function classes (Brownlees et al., 2015), but is computationally intractable due to the truncation and thus more of theoretical interest. In contrast, our ensemble approach is a meta algorithm that acts on any learning algorithm to provide exponential tail bounds regardless of the underlying problem characteristics. Relatedly, various techniques such as gradient clipping (Cutkosky & Mehta, 2021; Gorbunov et al., 2020) and MOM (Puchkin et al., 2024) have been adopted in stochastic gradient descent (SGD) algorithms for handling heavy-tailed gradient noises, but their focus is the faster convergence of SGD rather than generalization.
5 Conclusion
@@
This paper introduces a novel ensemble technique that significantly improves generalization by estimating the mode of the sampling distribution of the base learner via subsampling. In particular, our approach converts polynomially decaying generalization tails into exponential decay, thus providing order-of-magnitude improvements as opposed to constant factor improvements exhibited by variance reduction. Extensive numerical experiments in both machine learning and stochastic programming validate its effectiveness, especially for scenarios with heavy-tailed data and slow convergence rates. This work underscores the powerful potential of our new ensemble approach across a broad range of machine learning applications.
@@
The appendices are organized as follows. In Appendix A, we review additional related work. Appendix B presents additional technical discussion for Theorem 2.1. Next, in Appendix C, we document the proofs of the main theoretical results in our paper. Specifically, we introduce some preliminary definitions and lemmas in Appendix C.1. Then, the formal statement and the proof of Theorem 2.1 can be found in Appendix C.2. The formal statement and the proof of Theorem 2.3 can be found in Appendix C.4. To improve clarity, we defer the proofs for all technical lemmas to Appendix D. In Appendix E, we provide another motivating example that supplements Example 1.1. Finally, we provide additional numerical experiments in Appendix F.
@@ Appendix A Additional Related Work
Bagging has been adopted in stochastic optimization for various purposes. The most relevant line of works (Biggs et al., 2023; Perakis & Thayaparan, 2021; Wang et al., 2021; Biggs & Perakis, 2023) study mixed integer reformulations for stochastic optimization with bagging approximated objectives such as random forests and ensembles of neural networks with the ReLU activation. These works focus on computational tractability instead of generalization performance. Anderson & Nguyen (2020) empirically evaluates several statistical techniques including bagging against the plain SAA and finds bagging advantageous for portfolio optimization problems. Birge (2023) investigates a batch mean approach for continuous optimization that creates subsamples by dividing the data set into non-overlapping batches instead of resampling and aggregates SAA solutions on the subsamples via averaging, which is empirically demonstrated to reduce solution errors for constrained and high-dimensional problems. Another related batch of works (Lam & Qian, 2018a, b; Chen & Woodruff, 2024, 2023; Eichhorn & Römisch, 2007) concern the use of bagging for constructing confidence bounds for generalization errors of data-driven solutions, but they do not attempt to improve generalization. Related to bagging, bootstrap has been utilized to quantify algorithmic uncertainties for randomized algorithms such as randomized least-squares algorithms (Lopes et al., 2018), randomized Newton methods (Chen & Lopes, 2020), and stochastic gradient descent (Fang et al., 2018; Zhong et al., 2023), which is orthogonal to our focus on generalization performance.
Machine Learning for Optimization.
Learning to optimize (L2O) studies the use of machine learning in accelerating existing or discovering novel optimization algorithms. Much effort has been in training models via supervised or reinforcement learning to make critical algorithmic decisions such as cut selection (e.g., (Deza & Khalil, 2023; Tang et al., 2020)), search strategies (e.g., (Khalil et al., 2016; He et al., 2014; Scavuzzo et al., 2022)), scaling (Berthold & Hendel, 2021), and primal heuristics (Shen et al., 2021) in mixed-integer optimization, or even directly generate high-quality solutions (e.g., neural combinatorial optimization pioneered by (Bello et al., 2016)). See (Chen et al., 2022, 2024; Bengio et al., 2021; Zhang et al., 2023) for comprehensive surveys on L2O. This line of research is orthogonal to our goal, and L2O techniques can work as part of or directly serve as the base learning algorithm within our framework.
Appendix B Implications of Theorem 2.1 for Strong Base Learners
@@
We provide a brief discussion of Theorem 2.1 applied to fast convergent base learners. Based on Theorem 2.1, the way $p_{k}^{\max}$ and $\mathcal{E}_{k,\delta}$ enter into (10) reflects how the generalization performance of the base learning algorithm is inherited by our framework. To explain, large $p_{k}^{\max}$ and small $\mathcal{E}_{k,\delta}$ correspond to better generalization of the base learning algorithm. This can be exploited by the bound (10) with the presence of $\max\left\{1-p_{k}^{\max},\;\mathcal{E}_{k,\delta}\right\}$, which is captured with our sharper concentration of U-statistics with binary kernels. In particular, for base learning algorithms with fast generalization convergence, say $1-p_{k}^{\max}=\mathcal{O}(e^{-k})$ and $\mathcal{E}_{k,\delta}=\mathcal{O}(e^{-k})$ for simplicity, we have $C_{1}\max\left\{1-p_{k}^{\max},\;\mathcal{E}_{k,\delta}\right\}=\mathcal{O}(e^{-k})$ and hence the first term in (10) becomes $\mathcal{O}(e^{-n})$ which matches the error of the base learning algorithm applied directly to the full data set.
Appendix C Proofs for Main Theoretical Results
@@ C.1 Preliminaries
An important tool in the development of our theories is the U-statistic that naturally arises in subsampling without replacement. We first present the definition of U-statistic below and its concentration properties in Lemma C.2. The proof of Lemma C.2 can be found in Appendix D.1.
Definition C.1.
@@
Lemma C.2 (MGF dominance of U-statistics from (Hoeffding, 1963)).
For any integer $0<k\leq n$ and any kernel $\kappa(z_{1},\ldots,z_{k})$, let $U(z_{1},\ldots,z_{n})$ be the corresponding U-statistic defined in Definition C.1, and